# Bridge the Modality and Capacity Gaps in Vision-Language Model Selection

### v1

@@ 6 Conclusion
We analyze and address two key challenges in Language-Only VLM Selection (LOVM), which are VLM’s modality gap across different modal features and VLM’s capacity gap between its overall and dataset-specific rankings. Our key insight is that we can reuse the model’s useful statistics on open-source tasks to help the model selection on the target task. Swab utilizes a transport matrix between classes of the target task and open-source datasets to transfer VLM’s class-specific modality gap vectors and class-specific rank from open-source tasks to the target task, which mitigates the negative impacts of these two gaps. Experiment results on the LOVM benchmark show the superiority of our method.
@@ Appendix A LOVM Benchmark Details
LOVM Benchmark (Zohar et al., 2023) consists of 35 pre-trained VLMs and 23 datasets, with a total of $35\times 23=805$ evaluations. For each evaluation, LOVM provides the VLM’s zero-shot image classification accuracy on the corresponding dataset. Therefore, we can get the ground truth performance ranking of 35 VLMs on the 23 datasets.
@@ A.1 VLMs of LOVM Benchmark
To cover as many types of models as possible, the LOVM Benchmark uses OpenCLIP library (Ilharco et al., 2021) to get diverse VLMs. These VLMs differ from each other in terms of the model architecture (ResNet (He et al., 2016), Transformer (Vaswani et al., 2017), ConvNext (Liu et al., 2022)), the pre-trained dataset (OpenAI’s Data (Radford et al., 2021), LAION 2b (Schuhmann et al., 2021)), the training method (loss function/hyperparameter/data augmentation) and so on. Table 7 displays the relevant information of each VLM. The diversity of these VLMs ensures that the experimental results calculated on them can reflect the performance of the VLM model selection algorithm in real-world situations.
Table 7: The detailed information of 35 models used in the LOVM Benchmark. The table comes from  (Zohar et al., 2023).
@@ A.2 Datasets of LOVM Benchmark
Table 8: The detailed information of 23 tasks used in the LOVM Benchmark.
@@
Imagenet (Deng et al., 2009) 1000 classification natural image
SUN397 (Xiao et al., 2010) 397 scene und. natural image
Country211 (Radford et al., 2021) 211 geolocation natural image
Stanford Cars (Krause et al., 2013) 196 classification natural image
Flowers102 (Nilsback & Zisserman, 2008) 102 classification natural image
CIFAR100 (Krizhevsky & Hinton, 2009) 100 classification natural image
DTD (Cimpoi et al., 2014) 46 classification textural image
RESISC45 (Cheng et al., 2017) 45 classification satellite images
GTSRB (Stallkamp et al., 2011) 43 classification natural image
Oxford Pets (Parkhi et al., 2012) 37 classification natural image
VOC2007 (Everingham et al., 2007) 20 classification natural image
STL10 (Coates et al., 2011) 10 classification natural image
EuroSAT (Helber et al., 2019) 10 classification satellite images
MNIST (LeCun et al., 2010) 10 classification hand-writing
SVHN (Netzer et al., 2011) 10 OCR natural image
CLEVR-C (Johnson et al., 2017) 8 object counting natural image
CLEVR-D (Johnson et al., 2017) 8 distance est. natural image
FER2013 (Goodfellow et al., 2013) 7 fac. exp. rec. natural image
DMLab (Zhai et al., 2020) 6 distance est. synthetic
Retinopathy (Kaggle & EyePacs, 2015) 5 classification retina scan
KITTI (Geiger et al., 2013) 4 distance est. natural image
PCam (Veeling et al., 2018) 2 classification histopathology
Rendered SST2 (Radford et al., 2021) 2 OCR text image
@@ A.3 Evaluation Metrics of LOVM Benchmark
In LOVM, our aim is to maximize the rank similarity between the prediction of VLMs’ performance $\hat{\bm{r}}_{\mathcal{T}}=\{\hat{r}_{m,\mathcal{T}}\}_{m=1}^{M}$ and VLMs’ ground truth performance $\bm{r}_{\mathcal{T}}=\{r_{m,\mathcal{T}}\}_{m=1}^{M}$ on the target dataset, especially the rank similarity of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$. This is because we tend to focus only on whether the best models can be chosen. We use the following metrics to evaluate the rank similarity:
@@
  • Kendall’s Rank Correlation ($\tau$) – Kendall’s Rank Correlation $\tau$ measures the ranking consistency between two ranking lists. Here we mainly focus on whether models within the intersection of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$ have consistent rankings. Equation 23 shows the calculation method. The value of $\tau$ ranges from -1 to 1, indicating ranking correlation from complete inconsistency to complete consistency.
    $q_{ij}=\text{sign}(r_{i,\mathcal{T}}-r_{j,\mathcal{T}})\cdot\text{sign}(\hat{r}_{i,\mathcal{T}}-\hat{r}_{j,\mathcal{T}}),$ (22)
    $\tau=\begin{cases}0,&\text{if }|F|<2,\\ \textstyle\frac{2}{|F|(|F|-1)}\sum_{\begin{subarray}{c}i<j;i,j\in F\end{subarray}}q_{ij},&\text{otherwise}.\end{cases}$ (23)
@@ Appendix B ModelGPT Details
ModelGPT is a method proposed for LOVM (Zohar et al., 2023). In this section, we introduce the metrics that ModelGPT used in Equation 7.
@@ B.1 The Generation Process of Auxiliary Text Samples
ModelGPT (Zohar et al., 2023) utilizes ChatGPT (Ouyang et al., 2022) to generate auxiliary text data by designing prompts to query ChatGPT. This extra text data mainly includes the Captions Dataset and the Synonyms Dataset.
@@
ModelGPT uses the following prompt to guide LLM to generate realistic and confusing text data corresponding to the user-provided classes.
Generate long and confusing image captions for the {domain} domain, which will be used to evaluate a Vision-Language Model’s {task} performance. Generate 50 captions for {classname}:
We show some generated auxiliary text examples. For example, in the category of dog, one of the text samples generated by ChatGPT is ”An adorable dog perfect for cuddles and playtime.” ModelGPT collects the results from this prompt to form the captions dataset, $\bm{D^{\text{cap}}}$.
@@
Please list the superclasses/synonyms for {classname}. For example: chair: [furniture, seat, bench, armchair, sofa] {classname}:
@@ B.2 Text-Derived Scores
Fisher Criterion $\phi_{\text{fisher}}$. The Fisher score measures the closeness of these class prompt embeddings to one another. Equation 24 shows the calculation process of it where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $c_{i}$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes.
$\phi_{\text{fisher}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\theta(\hat{\bm{t}}_{i},\hat{\bm{t}}_{j})\right],$ (24)
Silhouette Score $\varphi_{\text{sil}}$. The Silhouette Score measures the separation of different-class samples in the caption dataset $\bm{D^{\text{cap}}}$. To calculate it, ModelGPT averages the cosine similarity of captions to the nearest other class by:
$\varphi_{\text{sil}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\frac{1}{N}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[j]_{k},\hat{\bm{t}}_{i})\right],$ (25)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{cap}}}[j]_{k}$ representing sample $k$ of class $j$ in the caption dataset $\bm{D^{\text{cap}}}$. There is a total of $N$ such samples for each class.
@@
$\rho_{\text{disp}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[i]_{k},\hat{\bm{t}}_{i}),$ (26)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in Radford et al. (2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{cap}}}[i]_{k}$ representing sample $k$ of class $i$ in the caption dataset $\bm{D^{\text{cap}}}$. There is a total of $N$ such samples for each class.
@@
$\gamma_{\text{syn}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{syn}}}[i]_{k},\hat{\bm{t}}_{i}),$ (27)
where $\hat{\bm{t}}_{i}$ is the class prompt embedding derived using the prompt ensemble strategies proposed in (Radford et al., 2021) for class $i$, $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{syn}}}[i]_{k}$ representing sample $k$ of class $i$ in the synonym dataset $\bm{D^{\text{syn}}}$. There is a total of $N$ such samples for each class.
Appendix C Implementation Details of Swab
In this section, we provide some details on the implementation of Swab, which are not mentioned in the main text due to space constraints, as well as details of some experiments related to Swab.
C.1 Filtering the Open-Source Tasks’ Classes
When the number of classes $|C_{\mathcal{S}}|$ in open-source datasets $\mathcal{S}$ is large, solving the optimal transport problem in Equation 11 can be time-consuming (as current optimal transport toolkits generally compute via CPU). To reduce the runtime of optimal transport, we can first filter the classes $C_{\mathcal{S}}$. Consider that only statistics of classes relevant to the target dataset are helpful. Therefore, we can filter out the classes in $C_{\mathcal{S}}$ that are irrelevant to the target dataset $\mathcal{T}$ based on the class-level textual semantic similarity between the open-source datasets’ classes and the target dataset’s classes. This process is shown in the following Equation:
@@ C.1 Filtering the Open-Source Tasks’ Classes
C.2 Using Partial Optimal Transport for Bridging the Capacity Gap
Partial optimal transport extends the optimal transport framework, enabling the selective transfer of elements from a source to a target distribution, rather than moving all elements. Its optimization problem is defined as in Equation 30. Here $z$ refers to the total amount of mass actually be transferred.
@@ C.2 Using Partial Optimal Transport for Bridging the Capacity Gap
$\quad\quad\bm{1}^{T}\bm{\gamma}^{T}\bm{1}=z\leq\min\left\{\|\bm{u}\|_{1},\|\bm{v}\|_{1}\right\}.$
We found that when using Equation 18 to bridge the Capacity Gap, the transport matrix $\bm{\gamma}^{*}$ obtained using partial optimal transport yields better results than the one obtained using the original optimal transport via solving the Equation 11. Therefore, in our implementation, we use the transport matrix derived from partial optimal transport to bridge the capacity gap. This also indicates that when estimating VLM’s statistics on the target dataset, different types of statistics have different preferences for the estimation methods used. This variability is worth further investigation.
Figure 5: Comparison of the consistency metrics between the accuracy calculated using text data before and after bridging the gap and the model’s true accuracy. After bridging the modality gap, the text data act as better image proxies to evaluate the model’s performance.
C.3 Data Normalization in Bridging the Modality Gap.
@@
Here $\bm{x}\in\mathbb{R}^{d}$ represents the image sample’s or text sample’s feature, while $\bm{\mu}\in\mathbb{R}^{d}$ and $\bm{\sigma}\in\mathbb{R}^{d}$ are calculated using the features of all samples of the same modality within its respective dataset. Besides, we also use z-score normalization to text data’s feature when we implement ModelGPT.
C.4 Task Groupings for Spider Chart Construction.
The task grouping in section 4.4 are as follows:
  • Medical: [Diabetic Retinopathy, PCam]
  • Remote: [EuroSAT, RESISC45]
  • Animals: [Oxford Pets]
  • Scene: [SUN397]
  • Traffic: [Stanford Cars, GTSRB, KITTI]
  • Universal: [ImageNet, STL10, VOC2007, CIFAR100]
C.5 Pseudo Code of Swab
Algorithm 1 shows the pseudo code of Swab.

### v3

@@ 5 Conclusion
We analyze and address two key challenges in Language-Only VLM Selection (LOVM), which are VLM’s modality gap across different modal features and VLM’s Capability gap between its overall and dataset-specific rankings. Our key insight is that we can reuse the model’s useful statistics on open-source datasets to help the model selection on the target dataset. Swab utilizes a transport matrix between classes of the target dataset and open-source datasets to transfer VLM’s class-specific modality gap vectors and class-specific rank from open-source datasets to the target dataset, which mitigates the negative impacts of these two gaps. Experiment results on the LOVM benchmark show the superiority of our method.
@@ Appendix A LOVM Benchmark Details
LOVM Benchmark [64] consists of 35 pre-trained VLMs and 23 datasets, with a total of $35\times 23=805$ evaluations. To further enhance the diversity of the VLM Zoo, we add some representative VLMs such as BLIP [28] and BEiT-3 [52] to expand the VLM Zoo. The final VLM Zoo contains a total of 43 models, with a total of $43\times 23=989$ evaluations. For each evaluation, LOVM provides the VLM’s zero-shot image classification accuracy on the corresponding dataset. Therefore, we can get the ground truth performance ranking of 43 VLMs on the 23 datasets.
@@ A.1 VLMs of LOVM Benchmark
To cover as many types of models as possible, the LOVM Benchmark uses OpenCLIP library [21] to get diverse VLMs. These VLMs differ from each other in terms of the model architecture (ResNet [17], Transformer [50], ConvNext [32]), the pre-trained dataset (OpenAI’s Data [43], LAION 2b [44]), the training method (loss function/hyperparameter/data augmentation) and so on. Table 7 displays the relevant information of each VLM. We further add BLIP and BEiT-3 to the original VLM Zoo. It should be noted that in addition to the image-text pair data listed in the Table 7, BEiT-3’s pre-training data also includes unimodal image dataset (ImageNet-21K) and text datasets (English Wikipedia, BookCorpus, OpenWebText, CC-News, Stories). The diversity of these VLMs ensures that the experimental results calculated on them can reflect the performance of the VLM model selection algorithm in real-world situations.
Table 7: The detailed information of 43 models used in the LOVM Benchmark. Some of the information in the table comes from  [64].
@@ 22 ViT-H-14 ViT-H/14 laion2b_s32b_b79k L2b-b
36 BLIP_retrieval_base_coco BLIP_retrieval_base_coco COCO+VG+CC+CC12M+SBU BLIP-Dataset
37 BLIP_retrieval_base_f30k BLIP_retrieval_base_f30k COCO+VG+CC+CC12M+SBU+Flickr30k BLIP-Dataset-f
38 BLIP_retrieval_large_coco BLIP_retrieval_large_coco COCO+VG+CC+CC12M+SBU BLIP-Dataset
39 BLIP_retrieval_large_f30k BLIP_retrieval_large_f30k COCO+VG+CC+CC12M+SBU+Flickr30k BLIP-Dataset-f
40 BEiT-3_retrieval_base_coco BEiT-3_retrieval_base_coco CC12M+CC3M+SBU+COCO+VG BEiT-Dataset
41 BEiT-3_retrieval_base_f30k BEiT-3_retrieval_base_f30k CC12M+CC3M+SBU+COCO+VG+Flickr30k BEiT-Dataset-f
42 BEiT-3_retrieval_large_coco BEiT-3_retrieval_large_coco CC12M+CC3M+SBU+COCO+VG BEiT-Dataset
43 BEiT-3_retrieval_large_f30k BEiT-3_retrieval_large_f30k CC12M+CC3M+SBU+COCO+VG+Flickr30k BEiT-Dataset-f
@@ A.2 Datasets of LOVM Benchmark
Table 8: Detailed information of 23 tasks used in the LOVM Benchmark. This table comes from [64].
@@
Imagenet [8] 1000 classification natural image
SUN397 [53] 397 scene und. natural image
Country211 [43] 211 geolocation natural image
Stanford Cars [25] 196 classification natural image
Flowers102 [39] 102 classification natural image
CIFAR100 [26] 100 classification natural image
DTD [5] 46 classification textural image
RESISC45 [4] 45 classification satellite images
GTSRB [47] 43 classification natural image
Oxford Pets [41] 37 classification natural image
VOC2007 [10] 20 classification natural image
STL10 [6] 10 classification natural image
EuroSAT [19] 10 classification satellite images
MNIST [27] 10 classification hand-writing
SVHN [38] 10 OCR natural image
CLEVR-C [23] 8 object counting natural image
CLEVR-D [23] 8 distance est. natural image
FER2013 [16] 7 fac. exp. rec. natural image
DMLab [62] 6 distance est. synthetic
Retinopathy [24] 5 classification retina scan
KITTI [15] 4 distance est. natural image
PCam [51] 2 classification histopathology
Rendered SST2 [43] 2 OCR text image
@@ A.3 Evaluation Metrics of LOVM Benchmark
In LOVM, our aim is to maximize the rank similarity between the prediction of VLMs’ ranking $\hat{\bm{r}}_{\mathcal{T}}=\{\hat{r}_{m}^{\mathcal{T}}\}_{m=1}^{M}$ and VLMs’ ground truth ranking $\bm{r}_{\mathcal{T}}=\{r_{m}^{\mathcal{T}}\}_{m=1}^{M}$ on the target dataset, especially the rank similarity of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$. This is because we tend to focus only on whether those appropriate models can be chosen. To ensure fair comparability, we follow the evaluation metrics used by LOVM [64] to assess different model selection methods and directly utilize the code provided by LOVM [64] to calculate these metrics:
@@
  • Kendall’s Rank Correlation ($\tau$) – Kendall’s Rank Correlation $\tau$ measures the ranking consistency between two ranking lists. We follow LOVM [64] to focus on the VLMs within the intersection of the top 5 VLMs in $\hat{\bm{r}}_{\mathcal{T}}$ and $\bm{r}_{\mathcal{T}}$ and use $\tau$ to evaluate a model selection method’s capability.
@@ Appendix B ModelGPT Details
ModelGPT is a method proposed by LOVM [64]. In this section, we introduce the metrics that ModelGPT used in Equation 3.
@@ B.1 The Generation Process of Auxiliary Text Samples
ModelGPT [64] utilizes ChatGPT [40] to generate auxiliary text data by designing prompts to query ChatGPT. This extra text data mainly includes the Captions Dataset and the Synonyms Dataset.
@@
ModelGPT uses the following prompt to guide LLM to generate realistic and confusing text data corresponding to the user-provided classes. The reason for requiring ChatGPT to generate confusing texts is to increase the classification difficulty of the text data, thereby enhancing its ability to distinguish the performance of different models.
Generate long and confusing image captions for the {domain} domain, which will be used to evaluate a Vision-Language Model’s {task} performance.
Generate 50 captions for {classname}:
We show some generated auxiliary text examples. For example, in the category of dog, one of the text samples generated by ChatGPT is "An adorable dog perfect for cuddles and playtime." ModelGPT collects the results from this prompt to form the captions dataset, $\bm{D^{\text{cap}}}$.
@@
Please list the superclasses/synonyms for {classname}. For example:
chair: [furniture, seat, bench, armchair, sofa]
{classname}:
@@ B.2 Text-Derived Scores
Fisher Criterion $\phi_{\text{fisher}}$. The Fisher score measures the closeness of VLM’s text classifier to one another. Equation 12 shows the calculation process of it where $\hat{\bm{t}}_{i}$ is the text classifier of the $i$-th class derived using the prompt ensemble strategies proposed in [43], $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes.
$\phi_{\text{fisher}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\theta(\hat{\bm{t}}_{i},\hat{\bm{t}}_{j})\right].$ (12)
Silhouette Score $\varphi_{\text{sil}}$. The Silhouette Score measures the separation of different-class samples in the caption dataset $\bm{D^{\text{cap}}}$. To calculate it, ModelGPT averages the cosine similarity of captions to the nearest other class’s classifier by:
$\varphi_{\text{sil}}=\frac{1}{|C|}\sum_{j=1}^{|C|}\text{max}_{i,i\neq j}\left[\frac{1}{N}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[j]_{k},\hat{\bm{t}}_{i})\right].$ (13)
where $\hat{\bm{t}}_{i}$ is the text classifier of the $i$-th class derived using the prompt ensemble strategies proposed in [43], $\theta(\cdot,\cdot)$ is a function that calculates the cosine similarity between two vectors, and $|C|$ is the number of classes. $\bm{D^{\text{cap}}}[j]_{k}$ representing sample $k$ of class $j$ in the caption dataset $\bm{D^{\text{cap}}}$. There is a total of $N$ such samples for each class.
@@
$\rho_{\text{disp}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{cap}}}[i]_{k},\hat{\bm{t}}_{i}).$ (14)
The definitions of all symbols in Equation 14 are consistent with those in Equation 13.
@@
$\gamma_{\text{syn}}=\frac{1}{|C|N}\sum_{i=1}^{|C|}\sum_{k=1}^{N}\theta(\bm{D^{\text{syn}}}[i]_{k},\hat{\bm{t}}_{i}).$ (15)
The definitions of $\hat{\bm{t}}_{i}$, $\theta(\cdot,\cdot)$, $|C|$ and $N$ in Equation 15 are consistent with those in Equation 13. $\bm{D^{\text{syn}}}[i]_{k}$ representing sample $k$ of class $i$ in the synonym dataset $\bm{D^{\text{syn}}}$.
Appendix C Related Work
Vision-Language Models. Vision-Language Models represent a class of multimodal models adept at correlating textual and visual information. These VLMs can comprehend rich semantics and possess strong generalization capabilities. Hence, they are often used in data-limited scenarios [54, 29]. Many VLMs are pre-trained or fine-tuned on extensive text-image pairs using loss functions such as contrastive loss, endowing them with powerful text-image matching capability. Prominent VLMs include CLIP [43], ALIGN [22], FLAVA [45], Florence [61], and CoCa [60]. These VLMs possess robust zero-shot image classification capabilities [43], which enables its widespread application in tasks characterized by long-tail distributions or those where collecting substantial training data is challenging, such as medical image analysis. Some work [37, 57] has also shown that by incorporating external knowledge, the zero-shot capabilities of CLIP can be further enhanced. In recent years, the number of open-source VLMs has been increasing [21]. Previous work [64] has pointed out that different VLMs possess varying image classification capabilities. This indicates that the performance of VLMs can vary significantly across different tasks and domains. These models with diverse capabilities constitute a VLM Zoo rich in knowledge. This VLM Zoo enables us to utilize different VLMs for various classification tasks, thereby changing the paradigm of using a single VLM to complete diverse classification tasks. This paper focuses on selecting the most suitable VLM for the target task from the VLM Zoo.
Model Selection.  The essence of the model selection problem lies in measuring the transferability of the model to the target task. Previous model selection methods [48, 49, 58, 59, 9, 20] evaluate the PTM’s transferability by performing a forward pass of the PTM on the target task’s data and calculating the metric of transferability based on the forward pass’s result. For example, H-Score [2], NCE [48], LEEP [49], LogME [58] estimate transferred log-likelihood, negative conditional entropy, log expectation, marginalized likelihood to obtain proxy metric of transferability, respectively. Some new methods, such as Task2Vec [1] and Model Spider [63], generate representation vectors for both the models and the tasks and measure the transferability of the model to the task by calculating the similarity between these vectors. However, VLMs are typically used in zero-shot or few-shot scenarios, where we are unable to obtain a large amount of data for the target task in advance, making traditional model selection approaches unsuitable for VLMs. Additionally, previous methods have mainly focused on single-modal models, overlooking the characteristics of VLMs. Therefore, in this paper, we concentrate on designing model selection algorithms that are suitable for scenarios with limited data and take into account the characteristics of VLMs.
Appendix D Implementation Details of Swab
In this section, we provide some details on the implementation of Swab, which are not mentioned in the main text due to space constraints.
D.1 Filtering the Open-Source Tasks’ Classes
The statistics of classes unrelated to the target class are generally not valuable for reuse. Meanwhile, when the number of classes $|C_{\mathcal{S}}|$ in open-source datasets $\mathcal{S}$ is large, solving the optimal transport problem in Equation 5 can be time-consuming (as current optimal transport toolkits generally compute via CPU). To reduce the runtime of optimal transport, we can first filter the classes $C_{\mathcal{S}}$. Consider that only statistics of classes relevant to the target dataset are helpful. Therefore, we can filter out the classes in $C_{\mathcal{S}}$ that are irrelevant to the target dataset $\mathcal{T}$ based on the class-level textual semantic similarity between the open-source datasets’ classes and the target dataset’s classes. This process is shown in the following Equation:
@@ D.1 Filtering the Open-Source Tasks’ Classes
D.2 Using Partial Optimal Transport for Bridging the Capability Gap
Partial optimal transport extends the optimal transport framework, enabling the selective transfer of elements from a source to a target distribution, rather than moving all elements. Its optimization problem is defined as in LABEL:pot_problem. Here $mass$ refers to the total amount of mass actually be transferred. We set $mass=0.9$ in our implementation.
@@ D.2 Using Partial Optimal Transport for Bridging the Capability Gap
$\quad\quad\bm{1}^{T}\bm{\gamma}^{T}\bm{1}=mass\leq\min\left\{\|\bm{u}\|_{1},\|\bm{v}\|_{1}\right\}.$
We found that when using Equation 8 to bridge the Capability Gap, the transport matrix $\bm{\gamma}^{*}$ obtained using partial optimal transport yields better results than the one obtained using the original optimal transport via solving the Equation 5. Therefore, in our implementation, we use the transport matrix derived from partial optimal transport to bridge the capability gap. This also indicates that when estimating VLM’s statistics on the target dataset, different types of statistics have different preferences for the estimation methods used. This variability is worth further investigation.
D.3 Data Normalization in Bridging the Modality Gap.
@@
Here $\bm{x}\in\mathbb{R}^{d}$ represents the image sample’s or text sample’s feature, while $\bm{\mu}\in\mathbb{R}^{d}$ and $\bm{\sigma}\in\mathbb{R}^{d}$ are calculated using the features of all samples of the same modality within its respective dataset.
D.4 Pseudo Code of Swab
Algorithm 1 shows the pseudo-code of Swab.